OpenStudy (anonymous):

Find the value of y for the given values of x and z y=x+z x=9, z=17

OpenStudy (anonymous):

y=x+z x=9, z=17 substitue y=9+17 y=?

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

welcome fan me and ill help u :)

OpenStudy (anonymous):

kk